The question is to write, and then solve this problem:
"The Length of a rectangle is 2 more than 8 times it's width. It's area is 36 cm^2. Find the width and length."
I wrote. (8w+2)w=36, and got the answers Length= 9\4, and width= 2. However when I look at the answer key, it says that the answer is Length= 18 and width = 2. I have no clue how they got this answer. PLEASE HELP IT IS DRIVING ME INSANE!!!

Answer by Alan3354(69443)

"The Length of a rectangle is 2 more than 8 times it's width. It's area is 36 cm^2. Find the width and length."
=====================
L = 8W + 2
L*W = 36
-----
W*(8W + 2) = 36
8W^2 + 2W - 36 = 0
4W^2 + W - 18 = 0
(W - 2)*(4W + 9) = 0
W = 2 --> L = 18
=================
The 2nd solution for W is -9/4, ignore it.
==================
PS it's = it is